OpenStudy (anonymous):

I don't remember how to do a problem like this. Can you please explain how to do it? Thank you!! Solve x + y + z = -6 x – y + z = -2 2x + y – 3z = 5 Infinite number of solutions no solutions (4, -6, -8) (-1, -2, -3)

OpenStudy (plasmataco):

the bottem

OpenStudy (plasmataco):

(-1,-2,-3)

OpenStudy (anonymous):

@Plasmataco can you help with this one too? Solve x + y + z = 12 2x – y – z = -6 x + 3y + 5z = 44